Kabbed:

Create your own piecewise function with at least two functions. Explain, using complete sentences, the steps for graphing the function. Graph the function by hand or using a graphing software of your choice (remember to submit the graph).

Kabbed:

@dude what is a piecewise function I need help

Kabbed:

@Shadow

dude:

Uh create 2 equations (one can be linear and the other could be parabolic)

Kabbed:

it says to write in complete sentences the steps what is the first step?

Kabbed:

this it the last question so I don't know if this question has to do with the other two

dude:

I am just trying to guide you, my goal is to get equations so you understand what to do and then we can format it to answer the question

Kabbed:

so how do I make a linear equation

Kabbed:

http://prntscr.com/mwfx07 is the slope in this linear equation positive or negative

Kabbed:

@dude

Kabbed:

@Aureo

Kabbed:

@umm @Pixel

Eiwoh2:

@AngeI

Kabbed:

@Elsa213 @Ultrilliam @pooja195 @KittyGirl

Kabbed:

@ZoeyBiocth16

Kabbed:

please even if you might not know even a little help is good

pooja195:

@AngeI @Vocaloid

BlankSpace:

I'm literally feeling for you so let's learn this together

BlankSpace:

You can practice and watch videos on Kahn academy. That's what I do

Kabbed:

sorry I was gone

Kabbed:

what's the first step to create a linear equation

Kabbed:

so a piecewise function is two functions together?

Kabbed:

@dude

Kabbed:

I don't know how to create a linear function

dude:

Oh hold on I've done this with someone before, let me refer you to that post

Kabbed:

Ok

dude:

https://questioncove.com/study#/updates/5c5883d760a4eae76bf54800 I would suggest to not use the same points, change it up a bit or else you might get caught plagiarizing

Kabbed:

y=6x-1 and y=x^4 @dude is that good? or should it be more different

dude:

Thats good

Kabbed:

I don't know what negative to -1 means

dude:

Hmm your first equation is \(6x-1\) or \(6x^{-1}\)?

Kabbed:

6x-1

dude:

Is it the first one or second one I wrote? That doesnt help xD

Kabbed:

the first one @dude

Kabbed:

https://prezi.com/njplm9bmowca/605-algebra-2-assignment/ this is how another student did the piecewise question

Kabbed:

@Shadow

Kabbed:

https://www.desmos.com/calculator

Kabbed:

hero help

Hero:

Are they expecting you to sketch the functions on paper or graphing it on a calculator?

Hero:

By the way, lines of the form y = mx + b are functions that are easy to graph piecewise.

Kabbed:

I can use desmos

Hero:

Piecewise functions are just two different functions that are GRAPHED on their own separate intervals along the x-axis.

Hero:

For example, I can graph \(f(x) = x + 2\) on the interval where \(x < 0\). I can then choose to graph \(g(x) = x^2\) on the interval \(x \ge 0\)

Hero:

Both those functions are easily graphed on desmos btw.

Kabbed:

I see @Hero sorry I'm back

Hero:

Have you attempted to graph the functions in desmos? If so, post a link to your work here.

Hero:

That link goes to a blank page.

Hero:

Here's the one of the easiest functions to graph: https://www.desmos.com/calculator/zchxzybgeh

Kabbed:

f(x)= x+1 x<2 and f(x) = x^4 x>0

Kabbed:

those are my two functions

Hero:

Except your piecewise function has to account for the entire real number domain from \((-\infty, \infty)\). You can split the piecewise function into as many intervals as you want, but it has to account for the entire domain.

Kabbed:

what

Hero:

Also there should not be any interval overlaps between each function piece.

Kabbed:

oh

Hero:

For what you are trying to graph, there is overlap between interval (0, 2).

Hero:

The format to graph a piecewise function in desmos is \(f(x) = \{\text{interval:function, interval:function}\}\)

Kabbed:

I can't do it

Hero:

For example, to graph \(y = x\) on interval \((0, \infty)\) and \(y = 1\) on interval \(-\infty, 0]\), you would write $$f(x) = \{x > 0: x, x \le 0: 1\}$$ https://www.desmos.com/calculator/7oe8mkwxi2

Kabbed:

what

Kabbed:

how do I do f(x)= x+1 x>2 f(x)= x + 2 x<0

Hero:

Is that the current problem you are working on?

Kabbed:

yes

Kabbed:

those are my two functions

Hero:

Well, I can't do it for you unfortunately.

Kabbed:

do what?

Kabbed:

\(f(x) = \{\text{x>2:x+1, x<0:x+2}\}\)

Hero:

Would you mind showing me the original problem your teacher gave you? Can you screenshot it on your computer or take a picture of it with your phone and upload it here?

Hero:

Piecewise functions do not have any overlap

Kabbed:

http://prntscr.com/mwjc3n

Kabbed:

f(x)= x+1 x>2 f(x)= x + 2 x<0 doesn't overlap

Hero:

Okay, there should not be any gaps either.

Hero:

Now there is a gap on \(0 < x < 2\)

Kabbed:

they both have to be on the same lines

Hero:

no laps or overlaps though.

Kabbed:

f(x)= x+1 x>0 f(x)= x + 2 x<0

Hero:

I can graph three functions: \(f(x) =\begin{cases} x^2, &x<0\\x+2, &0\le x <5\\\sin(x), &x \ge 5\end{cases}\)

Hero:

There is no gaps or overlaps. BTW, an open circle point is considered a gap.

Kabbed:

is https://www.desmos.com/calculator/hl34cgu3ia fine

Hero:

Yes but now a gap occurs at \(x = 0\)

Hero:

You've accounted for the interval between infinity and zero and between zero and infinity, but what happens when x = 0?

Kabbed:

I don't know

Hero:

you need to include the equal sign on one of those intervals

Hero:

Great effort so far though @Kabbed

Hero:

You're learning by trial and error. The best way to learn.

Kabbed:

f(x)= x+1 x<2 f(x)= x-4 x<3

Hero:

Have you tried graphing that in desmos?

Hero:

Your previous attempt was better. It was only missing an equal sign on one of the intervals. This current attempt is worse though because there is too much overlap on interval \(-\infty < x < 2\) and a big gap on interval \(3 < x < \infty\)

Kabbed:

how do I put a equal sign on one of them?

Hero:

add an equal sign next to the inequality sign.

Kabbed:

f(x)= x+1 x>=0 f(x)= x + 2 x<0

Hero:

Much better

Hero:

Now you just have to graph that in desmos.

Hero:

Yeah, that works, but it is the simplest kind of graph.

Kabbed:

x>0 is a constraint or a interval what do I call it

Hero:

In that form you can call it a constraint.

Hero:

Basically a constraint is a limitation you place on \(x\). You constrain it by saying the function is only valid when \(x > 0\). In other words, the function is only valid when \(x\) is greater than zero.

Hero:

If you write \((0, \infty)\) then you mean that the function exists on the interval between zero and infinity.

Hero:

But essentially \(x > 0\) and \((0,\infty)\) operate the same way. They both place limitations on the domain of the function.

Hero:

So if someone asks "What is the constraint on the function?". You can respond with something like "The constraint on the function is \(x > 0\).". However, if they ask "On what interval does the function exist?", you can say something like "The function exists on interval \((0, \infty)\)

Kabbed:

The next step, is to plot the rest of the points but only the ones to the left of 1 on the x-axis. Lastly connect the lines. Repeat the same steps for the second function but according to its numbers. @Hero

Kabbed:

ill try to write that^